# Internet of Mirrors for Connected Healthcare and Beauty: A Prospective Vision

### v1

@@ III-D Computer Vision
III-E Personalised Recommendation Systems
@@ III-G Computational Architectures
Table III: Overview of Enabling Technologies and Their Potential for IoM Integration
@@
Skin Imaging Methodologies
Classification and feature extraction, segmentation, capacitive imaging
Disease detection, wearable skin sensors
Early diagnosis, treatment recommendations
Real-Time Health Monitoring
PPG, Infrared thermography, RF Sensing
Vitals tracking, disease monitoring
Continuous monitoring, remote healthcare
Computer Vision
Augmentation of cosmetics
Makeup applications, retail scenarios
Real-time makeup previews, personalised shopping
Personalised Recommendation Systems
Deep learning, user feedback
Product suggestions, routine planning
User-specific recommendations, adaptive learning
Wireless Connectivity in IoT
5G, LPWAN, WLAN, Tactile Internet
Wireless data transfer, remote control
Low latency, high data rates, scalability
Computational Architectures
Edge Computing, Cloud Computing, Distributed Computing
Data Processing, Scalable Resources
Real-time monitoring, Large dataset handling
IV IoM Scenarios and Use Cases
We expect that the realisation of the IoM will allow many everyday activities and aspects of human life to be re-imagined with respect to a technological ecosystem of interactions between consumers (customers, patients etc.), professionals (dermatologists, makeup artists etc.) and service providers (businesses, retail stores etc.). To present an idea and visualisation of what these interactions could be like and what the implications would be, five hypothetical scenarios that could potentially arise from the realisation of the IoM are presented below as well as the corresponding missing capabilities and considerations. Following on from Fig 4, a visualisation of these scenarios with the possible interactions is presented in Fig 6.
Figure 6: Visualisation of IoM scenarios and possible interactions
IV-A Scenario 1 (Enhanced Beauty Consultations):
@@
IV-B Scenario 2 (Employee Training):
@@
IV-C Scenario 3 (Eco-Friendly Inventory Management):
@@
IV-D Scenario 4 (Remote Dermatological Treatment and Skin Health Analysis):
@@
IV-E Scenario 5 (All-Inclusive Remote Living Support for the Smart Home):
@@
V Current Challenges and Future Research Directions
This section addresses the challenges associated with the development and realisation of the IoM. It highlights key obstacles including communication protocol performance, system interoperability, multimodal content representation and data privacy and security. Possible solutions to these challenges that can support the practical realisation of IoM systems have also been presented.
V-A Low Latency and High Reliability
@@ V-B Interoperability and Standardisation
V-C Representation and Analysis of Multimodal Content
@@
V-D Privacy and Security Issues

### v2

@@ 3.4 Computer Vision
3.5 Personalised Recommendation Systems
@@ 3.7 Computational Architectures
Table 3: Overview of Enabling Technologies and Their Potential for IoM Integration
@@
Skin Imaging Methodologies Classification and feature extraction, segmentation, capacitive imaging Disease detection, wearable skin sensors Early diagnosis, treatment recommendations
Real-Time Health Monitoring PPG, Infrared thermography, RF Sensing Vitals tracking, disease monitoring Continuous monitoring, remote healthcare
Computer Vision Augmentation of cosmetics Makeup applications, retail scenarios Real-time makeup previews, personalised shopping
Personalised Recommendation Systems Deep learning, user feedback Product suggestions, routine planning User-specific recommendations, adaptive learning
Wireless Connectivity in IoT 5G, LPWAN, WLAN, Tactile Internet Wireless data transfer, remote control Low latency, high data rates, scalability
Computational Architectures Edge Computing, Cloud Computing, Distributed Computing Data Processing, Scalable Resources Real-time monitoring, Large dataset handling
4 IoM Scenarios and Use Cases
We expect that the realisation of the IoM will allow many everyday activities and aspects of human life to be re-imagined with respect to a technological ecosystem of interactions between stakeholders within the IoM ecosystem. These stakeholders includes:
  • 1.
    Consumers: Individuals such as customers and patients who interact with the IoM system to receive personalised services and recommendations.
  • 2.
    Professionals: Specialists like dermatologists, makeup artists, and other professionals who provide expertise and services through the IoM platform.
  • 3.
    Service Providers: Entities such as businesses, retail stores, and healthcare providers that offer products and services integrated into the IoM system.
To present an idea and visualisation of what these interactions could be like and what the implications would be, five hypothetical use case scenarios that could potentially arise from the realisation of the IoM are presented below as well as the corresponding missing capabilities and considerations. Following on from Fig 4, a visualisation of these use case scenarios with all the relevant stakeholders and the possible interactions between them all within the ecosystem of the IoM are presented in Fig 6
Figure 6: Visualisation of IoM Use Cases and Possible Stakeholder Interactions
4.1 Scenario 1 (Enhanced Beauty Consultations):
@@
4.2 Scenario 2 (Employee Training):
@@
4.3 Scenario 3 (Eco-Friendly Inventory Management):
@@
4.4 Scenario 4 (Remote Dermatological Treatment and Skin Health Analysis):
@@
4.5 Scenario 5 (All-Inclusive Remote Living Support for the Smart Home):
@@
Unique Advantages
While the scenarios presented above may share similarities with existing technologies, the IoM offers unique advantages that set it apart. The IoM is not designed to replace smartphones, tablets, or other smart devices, but rather to complement them by providing a specialised, visual interface that seamlessly integrates into daily life, particularly in health and beauty contexts to offer comprehensive monitoring, real-time full-body analysis, and personalised recommendations through an always-on display that enhances user experience beyond what handheld devices can provide.
The key differentiators of the IoM include its passive integration into daily routines, requiring minimal active engagement from users, and its role as a unified platform for health and beauty services, ensuring data consistency and interoperability. Unlike fragmented apps and services, the IoM provides a centralised ecosystem with a large, always-on visual interface that offers full-body visualisation and interaction, which is particularly suited for health and beauty applications. This design allows for hands-free operation, making it ideal for scenarios where manual interaction with devices is impractical.
Furthermore, the IoM can facilitate more standardised and seamless integration of professional services, from beauty consultations to healthcare diagnostics. Its fixed position allows for consistent, long-term monitoring of user health and appearance, enabling more accurate tracking of changes over time. These features position the IoM as a novel, unified platform that addresses specific needs in the health and beauty domains, offering capabilities that go beyond what current fragmented solutions can provide. Table 4 provides a more comprehensive comparison between the IoM and existing systems and outlines the specific advantages of the IoM and how they address limitations in current devices.
Table 4: Detailed Breakdown of IoM’s Unique Advantages
Feature Description Advantage
Full-body, hands-free analysis Comprehensive full-body scans and virtual try-ons without handheld devices. Overcomes limitations of smaller, handheld devices for full-body assessments.
Passive, continuous monitoring Monitors health and beauty metrics over extended periods without active user intervention. Complements wearables, offering continuous monitoring without charging or wearing devices.
Multi-user, shared experience Enables multi-user interactions and community-based features in shared spaces. Facilitates shared experiences not easily achieved with personal devices.
Professional-grade imaging Consistent lighting and high-resolution cameras for detailed skin analysis. Surpasses varying quality of phone cameras and lighting conditions.
Contextual recommendations Provides environment-aware health and beauty recommendations. Offers more consistent environmental tracking than mobile devices.
Seamless professional integration Standardised platform for remote professional consultations. Provides more consistent, high-quality visual data than varied personal devices.
These unique capabilities position the IoM as a complementary technology to existing devices, filling crucial gaps in current health and beauty tech ecosystems. By working alongside and enhancing the capabilities of smartphones, tablets, and wearables, the IoM provides a more comprehensive and integrated approach to personal health and beauty management.
@@
5 Current Challenges and Future Research Directions
This section addresses the challenges associated with the development and realisation of the IoM. It highlights key obstacles including communication protocol performance, system interoperability, multimodal content representation and data privacy and security. Possible solutions to these challenges that can support the practical realisation of IoM systems have also been presented. .
5.1 Low Latency and High Reliability
  • 1.
@@ 5.2 Interoperability and Standardisation
    In addition, leveraging industry standards and best practices can also enhance the development and adoption of the IoM framework. For example, the TM Forum [161], a global industry association which provides a comprehensive suite of frameworks and best practices that can be particularly valuable in this context. The TM Forum’s frameworks, such as the Business Process Framework (eTOM) and the Information Framework (SID), offer structured methodologies for managing business processes and information systems in a scalable and efficient manner. Integrating these frameworks into the IoM concept can ensure better alignment with industry standards, enhance interoperability, and provide a robust foundation for the future development of the IoM ecosystem. Exploring and adopting relevant aspects of the TM Forum’s frameworks can thus contribute to overcoming some of the challenges identified and support the realisation of a more resilient and standardised IoM infrastructure.
5.3 Representation and Analysis of Multimodal Content
  • 1.
@@ 5.3 Representation and Analysis of Multimodal Content
5.4 Privacy and Security Issues